OpenStudy (anonymous):

Two lines are related by the scale 4 millimeters: 5 decimeters. how many times larger is the larger line?

OpenStudy (anonymous):

someone please help me

OpenStudy (anonymous):

thanks!

OpenStudy (skullpatrol):

milli = 1/(10^3) deci = 1/10

OpenStudy (skullpatrol):

4 millimeters = 0.04 decimeter

OpenStudy (skullpatrol):

5/(0.04)

OpenStudy (anonymous):

125?